# Leveraging Peer, Self, and Teacher Assessments for Generative AI-Enhanced Feedback

### v1

@@ VI Conclusions and Future Work
Future work will focus on validating this enhanced model through classroom studies, examining its impact on feedback quality, student engagement, and calibration of self-assessment skills. In this regard, it will be especially relevant to investigate how different forms of self-assessment guidance interact with GenAI-enhanced feedback. Prior research has shown that analytic rubrics can increase performance but may also reduce learning-oriented self-regulation, whereas self-assessment scripts tend to foster metacognitive monitoring and process-focused reflection, even if they offer less explicit guidance on expected performance standards [panadero2014rubrics]. Building on these findings, future iterations of AICoFe could experiment with alternative or hybrid scaffolds (e.g., rubric-only, script-based, or dynamically adapted combinations) and examine their differential effects on students’ self-regulation, perceived workload, and bias patterns in peer and self-assessment.
In addition, we plan to extend AICoFe’s data integration layer by incorporating multimodal learning analytics, including biosensor [becerra2025enhancing_review] and behavioral data, to capture affective and physiological dimensions of learning [becerra2025ai].

### v2

@@ VI Conclusions and Future Work
Future work will focus on validating this enhanced model through classroom studies, examining its impact on feedback quality, student engagement, and calibration of self-assessment skills. In this regard, it will be especially relevant to investigate how different forms of self-assessment guidance interact with GenAI-enhanced feedback. Prior research has shown that analytic rubrics can increase performance but may also reduce learning-oriented self-regulation, whereas self-assessment scripts tend to foster metacognitive monitoring and process-focused reflection, even if they offer less explicit guidance on expected performance standards [25]. Building on these findings, future iterations of AICoFe could experiment with alternative or hybrid scaffolds (e.g., rubric-only, script-based, or dynamically adapted combinations) and examine their differential effects on students’ self-regulation, perceived workload, and bias patterns in peer and self-assessment.
In addition, we plan to extend AICoFe’s data integration layer by incorporating multimodal learning analytics [2], including biosensor [4] and behavioral data, to capture affective and physiological dimensions of learning [7].
@@ Acknowledgment
Support by projects: Cátedra ENIA UAM-VERIDAS en IA Responsable (NextGenerationEU PRTR TSI-100927-2023-2), M2RAI (PID2024-160053OB-I00, MICIU/FEDER) and SNOLA (RED2022-134284-T). Alvaro Becerra is funded by a predoctoral contract (FPI) from the Comunidad de Madrid (PIPF-2024/COM-34288).